A sodium sulfate solution was prepared by dissolving 825 g of Sodium sulfate in 1850 ml...

  1. A sodium sulfate solution was prepared by dissolving 825 g of Sodium sulfate in 1850 ml of water (H 2 O). It was heated at 21.0 . C to 75.5 . C to dissolve the sodium sulfate in water. Specific gravity of sodium sulfate is 0.95 Cal /g . C. Calculate to determine the total energy used in KJ for making the solution.
Answer #1

Heat absorbed by the sodium sulfate Solution(q) = m*s*DT

m = mass of sodium sulfate solution taken = 825+1850 = 2675 g ( density of water = 1 g/cc)

s = specific heat of sodium sulfate solution = 0.95 cal/g.c

DT = 75.5 - 21.0 = 54.5 C

q = 2675*0.95*54.5 = 138.5 kcal

= 138.5*4.184 kj

= 579.5 kj